OpenStudy (anonymous):

I GIVE MEDALS AND FAN! Write the equation of the given circle. center (3, 3) radius of 2

OpenStudy (anonymous):

genreal eq is x2+y2=c2 here x=y=3 and c=2 put values and get result...

OpenStudy (anonymous):

I'm confused. Is x2+y2=c2 the answer?

OpenStudy (anonymous):

Cause it's supposed to be _x______y______

OpenStudy (anonymous):

@mathstudent55 @rockstar0765

OpenStudy (anonymous):

so it is.. (x-x1)^2+(y-y1)^2=c^2 x1=y1=3 c=2